### v2

# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

Abstract
The discovery of high-temperature superconductivity (SC) (HTSC) in pressurized La3Ni2O7 with critical temperature $T_{c}$ higher than the boiling point of liquid nitrogen has aroused a surge in the exploration of HTSC in the Ruddlesden-Popper phase multilayer nickelates. Very recently, SC is found in the La3Ni2O7 ultrathin film grown on the SrLaAlO4 substrate with $T_{c}$ above the McMillan limit ($\approx 40\text{ K}$) at ambient pressure (AP), allowing various experimental investigation on the pairing mechanism in this material. It is now eager to enhance the $T_{c}$ of La3Ni2O7 at AP. Here we propose that an imposed strong perpendicular electric field can strongly enhance the $T_{c}$ in the single-bilayer film of La3Ni2O7 at AP. The physics underlying this proposal is clear and simple. Under strong electric field, the layer with lower potential energy will accept electrons flowing from the other layer to fill in the Ni-$3d_{x^{2}-y^{2}}$ orbitals in this layer, as the nearly half-filled Ni-$3d_{z^{2}}$ orbital in this layer cannot accommodate more electrons. With the enhancement of the filling fraction in the $3d_{x^{2}-y^{2}}$ orbitals in this layer, the interlayer $s$-wave pairing will be subjected to the pair-breaking effect and be suppressed, but the intralayer $d$-wave pairing in this layer is promptly and strongly enhanced, which mimics the cuprates. Our combined simplified one-orbital study and comprehensive two-orbital one under the mean-field treatment and the density matrix renormalization group approach consistently verify this idea and yield that an imposed voltage of about $0.1\sim 0.2$ volt between the two layers is enough to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in this single bilayer at AP. Our results appeal for experimental verification.
The discovery of superconductivity (SC) with critical temperature $T_{c}$ above the boiling point of liquid nitrogen ($\approx 77\text{ K}$) in the pressurized La3Ni2O7 [1, 2, 3, 4, 5, 6, 7, 8, 9] has attracted great interests [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144]. This discovery has sparked the exploration of high-temperature SC (HTSC) in Ruddlesden-Popper phase multilayer nickelates, resulting in the discovery of SC in the pressurized La4Ni3O10 [39, 40, 41, 42, 43, 44, 45], which together with the previously synthesized infinite-layer nickelates Nd1-xSrxNiO2 [46, 47, 48, 49] have established a new family of SC other than cuprates and iron-based superconductors. However, the high pressure (HP) circumstance not only strongly hinders the experimental detection of the samples but also brings difficulties in the application of SC in industry. Very recently, the La3Ni2O7 ultrathin film with a few layers of unit cell grown on the SrLaAlO4 (SLAO) substrate has been grown by two different groups independently and SC with $T_{c}$ above the McMillan limit ($\approx$40 K) has been detected at ambient pressure (AP)[132, 133, 134], allowing various experimental investigation of the pairing mechanism in this material, attracting a lot of interests [135, 136, 137, 138, 139, 140, 141, 142, 143, 144]. It is now eager to enhance the $T_{c}$ of this material at AP. Here we propose a viable approach to realize $T_{c}$ above the boiling point of liquid nitrogen in the La3Ni2O7 single-bilayer film at AP.
Presently, the pairing mechanism in the La3Ni2O7, either in the bulk material under HP [74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 104, 105, 106, 107, 108, 125, 126, 127, 128] or in the ultrathin film at AP [135, 138, 140, 143], is still under debate. Density-functional-theory (DFT) based first-principle calculations have suggested that the low-energy orbitals are mainly Ni-$3d_{z^{2}}$ and $3d_{x^{2}-y^{2}}$, which are nearly half- and quarter- filled [50, 51, 52, 53, 54, 55, 56, 57, 58, 61]. Various experiments have revealed the strongly-correlated characteristic of the material [14, 16, 22, 24, 25, 28, 31, 32]. Particularly, the optical study reveals significant reduction of the electron kinetic energy which places the system in the proximity of the Mott phase [25]; the angle-resolved photoemission spectroscopy uncovers strong band renormalization caused by electron correlation [28]; the linearly temperature-dependent resistivity suggests “strange-metal” behavior [2]. Therefore, we take a strong-coupling view of the system. Under the strong Hubbard repulsion, the nearly half-filled $3d_{z^{2}}$ electrons can almost be viewed as localized spins. Therefore, the main carrier of SC should be the $3d_{x^{2}-y^{2}}$ electrons, which subject to the in-plane superexchange interaction just mimic the 50% hole-doped cuprates. However, it is a problem how HTSC can emerge under such a high doping level. The key physics lies in the important role played by the $3d_{z^{2}}$ orbitals. Through strong interlayer superexchange, the $3d_{z^{2}}$ electrons form interlayer pairing. The interlayer superexchange or the interlayer pairing of the $3d_{z^{2}}$ electrons can be transmitted to the $3d_{x^{2}-y^{2}}$ electrons through the Hund’s rule [79, 80, 82, 87, 92, 93, 95, 96, 97, 98, 100, 107, 108] or the nearest-neighbor (NN) hybridization [65, 83, 86, 87, 91, 98, 105, 107, 108] or both. Under such view, the role of pressure in enhancing the $T_{c}$ lies in the enhancement of the interlayer superexchange, the inter-orbital hybridization, or both.
Figure 1: (a) Schematic diagram for the dominant hopping integrals and superexchange interactions between the $E_{g}$ orbitals in La3Ni2O7. (b) Schematic diagram illustrating that the Hund’s rule coupling transmits the interlayer superexchange interaction $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the effective one $\tilde{J}_{\perp}$ between the $3d_{x^{2}-y^{2}}$ orbitals.
In this work, we propose an alternative approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the ultrathin film of La3Ni2O7 at AP. Here we consider the thinnest limit, i.e. a single bilayer film of La3Ni2O7. We can impose a perpendicular electric field, say pointing upward, in this single bilayer, so that electrons from the top layer will flow to the bottom layer. These electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. The enhancement of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number will first suppress the interlayer $s$-wave SC due to mismatch of the electron numbers between the two layers, and then promptly lead to the intra-bottom-layer $d$-wave SC with strongly enhanced $T_{c}$. To test this idea, we have performed a combined simplified single orbital study and a comprehensive two orbital study, which consistently yield that a voltage of experimentally achievable levels (around $0.1\sim 0.2$ volt predicted by the mean-field calculations) between the two layers is enough to induce $d$-wave SC with $T_{c}$ above the boiling point of liquid nitrogen in the bottom layer. Intriguingly, the $d$-wave SC carried by the bottom layer $3d_{x^{2}-y^{2}}$ electrons coexists with the interlayer $s$-wave pseudo gap carried by the $3d_{z^{2}}$ electrons in the mixing ratio of $1:\mathrm{i}$, breaking time-reversal symmetry. Our proposal potentially provides a viable approach to realize HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single bilayer film of La3Ni2O7.
I Results
I.1 Consideration and a Simplified Study
The La3Ni2O7 ultrathin film grown on the SLAO substrate form a bilayer square lattice [137, 135]. As illustrated in Fig. 1 (a), the leading hopping integrals are the interlayer hopping of the $3d_{z^{2}}$ electrons $t_{\perp}$ and the intralayer NN hopping of the $3d_{x^{2}-y^{2}}$ electrons $t_{\parallel}$. Under strong Hubbard $U$, these hopping terms can induce the effective superexchange $J_{\perp}$ and $J_{\parallel}$ through $J\approx\frac{4t^{2}}{U}$. Under the Hund’s rule coupling $J_{H}$, the spins of the two orbitals are inclined to be parallel aligned, as illustrated in Fig. 1 (b), which partly transmits the interlayer superexchange $J_{\perp}$ between the $3d_{z^{2}}$ orbitals to the $3d_{x^{2}-y^{2}}$ orbitals as $\tilde{J}_{\perp}=\alpha J_{\perp}$ with $\alpha\in(0,1)$. In addition, there exists intralayer NN- hybridization $t_{xz}$ between the two orbitals. As shown in Fig. 2 (a, b), the nearly quarter-filled $3d_{x^{2}-y^{2}}$ electrons subject to $J_{\parallel}$ and $\tilde{J}_{\perp}$ form interlayer-dominant pairing [79].
Figure 2: (a) Particle numbers of the four $E_{g}$ orbitals within an unit cell without electric field. (b) The dominant pairing configuration for (a). (c) Schematic diagram showing how the electrons flow under the perpendicular electric field $\bm{\varepsilon}$ pointing upward. (d) The dominant pairing configuration for (c).
Now let us turn on the upward electric field $\bm{\varepsilon}$, forcing electrons downward, as shown in Fig. 2 (c, d). In the top layer, since the $3d_{z^{2}}$ orbitals host larger density of state (DOS) than the $3d_{x^{2}-y^{2}}$ orbitals, they will donate more electrons. Most of these donated electrons will fill the $3d_{x^{2}-y^{2}}$ orbitals in the bottom layer, as the nearly half-filled $3d_{z^{2}}$ orbitals there cannot accommodate more electrons. A minority of the donated electrons can also be accepted by the top-layer $3d_{x^{2}-y^{2}}$ orbitals.
Figure 3: (a) The pairing amplitude $\tilde{\Delta}$ (in unit of $t_{\parallel}$) as function of the bottom-layer particle number per site $n_{bx}$. Different pairing symmetries are distinguished by color. (b) The $T_{c}$ as function of $n_{bx}$, in comparison with $0.42\tilde{\Delta}$ for the $d$-wave and $0.9\tilde{\Delta}$ for the $s$-wave regime. Inset: the spinon pairing temperature $T_{\mathrm{pair}}$ and the holon condensation temperature $T_{\mathrm{BEC}}$ as function of $n_{bx}$. In (a,b), we set $J_{\parallel}=0.4t_{\parallel}$ and $\tilde{J}_{\perp}=(1-\delta_{tz})\times 1.3J_{\parallel}$. (c)-(d) The pairing configurations of the $s$-wave and $d$-wave, respectively.
Even with doped holes under $\bm{\varepsilon}$, the top-layer $3d_{z^{2}}$ electrons cannot carry SC: Firstly, lacking pairing interaction, they cannot form intralayer pairing. Secondly, although they can pair with the localized bottom-layer $3d_{z^{2}}$ electrons, such pairs cannot coherently move, only resulting in the pseudo-gap. Therefore, the SC under $\bm{\varepsilon}$ can only be carried by the $3d_{x^{2}-y^{2}}$ orbitals. As the filling fractions of the $3d_{x^{2}-y^{2}}$ orbitals in the two layers are different, their Fermi levels are relatively shift, leading to mismatch of their Fermi surfaces (FSs), which will suppress their interlayer pairing. The bottom-layer $3d_{x^{2}-y^{2}}$ orbitals will form $d$-wave SC, mimicsing the cuprates, as shown in Fig. 2 (d). When $\bm{\varepsilon}$ is strong enough so that the filling fraction of the bottom-layer $3d_{x^{2}-y^{2}}$ orbitals is near that of the optimal doped cuprates, $d$-wave HTSC with strongly enhanced $T_{c}$ will be achieved in this layer, and the top layer will also acquire SC below $T_{c}$ through proximity.
Based on the above general consideration, we first conduct the following simplified model study including only the $3d_{x^{2}-y^{2}}$-orbital, with the $3d_{z^{2}}$ orbital only viewed as a source which tunes the total electron number. The following widely adopted single $3d_{x^{2}-y^{2}}$-orbital bilayer $t-J-J_{\perp}$ model [79, 80, 82, 92, 93, 95, 96, 100] is adopted,
$H=$displaystyle-- $t_{\parallel}\sum_{\left\langle{}i,j\right\rangle,\mu,\sigma}\left(c^{{\dagger}}_{i\mu\sigma}c_{j\mu\sigma}+\mathrm{h.c.}\right)+\sum_{i,\mu}\epsilon_{\mu}n_{i\mu}$ (1)
$+$ $J_{\parallel}\sum_{\left\langle{}i,j\right\rangle,\mu}\bm{S}_{i\mu}\cdot\bm{S}_{j\mu}+\tilde{J}_{\perp}\sum_{i}\bm{S}_{it}\cdot\bm{S}_{ib}.$
Here $c^{{\dagger}}_{i\mu\sigma}$ creates an electron at site $i$ in the layer $\mu$ (=top ($t$)/bottom ($b$)) with spin $\sigma$, and $n_{i\mu}$ or $\bm{S}_{i\mu}$ denote the corresponding electron number or spin operator. Only NN- bond $\langle i,j\rangle$ is considered in the summation. The $\epsilon_{\mu}$ is introduced to control the filling fractions of the two layers under $\bm{\varepsilon}$, with $\epsilon_{t}-\epsilon_{b}=\varepsilon$. However, as the total particle number of the $d_{x^{2}-y^{2}}$ electrons under given $\bm{\varepsilon}$ is unknown, we have to assume the ratio $r:1$ between the electron number flowing from the $3d_{z^{2}}$ orbitals and that flowing from $3d_{x^{2}-y^{2}}$ orbitals in the top layer when solving the model with the standard slave-boson mean-field (SBMF) theory [145]. Due to reason of DOS, we assume this ratio to be $2:1$, with details provided in the Supplementary Information (SI) [146]. Nevertheless, the concrete value of this ratio turns out not to obviously affect the results (see the SI [146]). The filling fractions are fixed under this assumption in the SBMF study. To capture the quantum fluctuation beyond mean-field, the density matrix renormalization group (DMRG) [147, 148] method is also employed, whose results are qualitatively consistent with those of the SBMF study. See details in the METHODS and SI [146].
We set $t_{\parallel}=1$ as the energy unit and $J_{\parallel}=0.4t_{\parallel}$ in our study. $\tilde{J}_{\perp}=(1-\delta_{tz})\times 1.3J_{\parallel}$ is applied in our SBMF study. The results are shown in Fig. 3. Fig. 3 (a) shows the amplitude and symmetry of the ground-state pairing gap as function of the bottom-layer $3d_{x^{2}-y^{2}}$ electron number $n_{bx}$, whose value enhances with $\bm{\varepsilon}$. It is shown that when $\bm{\varepsilon}$ or $n_{bx}$ enhances, the pairing amplitude $\tilde{\Delta}$ decays first and then increases. When $n_{bx}=0.5$, the ground state is interlayer $s$-wave SC. When $n_{bx}\geq 0.53$, the ground state is an intralayer $d$-wave SC, with the dominant pairing limited in the bottom layer. It is inspiring that with the enhancement of $n_{bx}$ in this regime, the $\tilde{\Delta}$ enhances promptly, similar to the case in the overdoped cuprates, wherein the enhancement of the filling fraction promptly enhances the pairing strength. The pairing configurations of the two different pairing symmetries are illustrated in Fig. 3 (c-d).
The $T_{c}\sim n_{bx}$ is shown in Fig. 3 (b). In the SBMF theory, the $T_{c}$ is given as the lower one between the spinon-pairing temperature $T_{\text{pair}}$ and the holon-BEC temperature $T_{\text{BEC}}$. The inset of Fig. 3 (b) displays $T_{\text{BEC}}\gg T_{\text{pair}}$, rendering $T_{c}=T_{\text{pair}}$ in the considered $n_{bx}$ regime. Note that the $T_{c}$ here is in the sense of Kosterlitz-Thouless transition. A comparison between Fig. 3 (b) and (a) suggests that $T_{c}$ scales with $\tilde{\Delta}$, which is more clear when the $T_{c}\sim n_{bx}$ is well fitted by $0.42\tilde{\Delta}\sim n_{bx}$ for the $d$-wave and $0.9\tilde{\Delta}\sim n_{bx}$ for the $s$-wave in Fig. 3 (b), consistent with the Bardeen-Cooper-Schrieffer (BCS) theory. Inspiringly, for $n_{bx}\geq 0.75$, the $T_{c}\gtrsim 0.02t_{\parallel}\approx 80$ K, suggesting the HTSC in the liquid nitrogen temperature range.
On the above, we have adopted $\tilde{J}_{\perp}=\alpha J_{\perp}(1-\delta_{tz})$ with $\alpha=1$, where $\delta_{tz}$ denotes the hole density of the top-$3d_{z^{2}}$ orbital. For the reduced $\alpha$, only the low-$n_{bx}$ regime accommodating the interlayer $s$-wave pairing in Fig. 3 (a, b) shrinks but the high-$n_{bx}$ regime accommodating the intralayer $d$-wave SC is not affected because the intralayer pairing is blind to $\tilde{J}_{\perp}$. Furthermore, assuming different ratios between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals turns out to yield similar results when expressed as functions of $n_{bx}$, as the dominant pairing under strong $\bm{\varepsilon}$ is the intra-bottom-layer pairing, which is blind to the filling fraction of the top layer. See the SI for details [146].
We have further employed the DMRG approach, which can capture the quantum fluctuation beyond mean-field, to compute the ground state of Hamiltonian (1) under different electric fields $\varepsilon$ and the electron-doping levels of the $d_{x^{2}-y^{2}}$ orbitals $\delta=n_{tx}+n_{bx}-1$. For $\varepsilon=0$, we have $\delta=0$. When $\varepsilon$ increases, it drives electrons from $d_{z^{2}}$-orbitals in the top layer to $d_{x^{2}-y^{2}}$-orbitals in both layers, increasing $\delta$. However, as the exact relationship between $\varepsilon$ and $\delta$ is unclear, we set them as two independent variables in our DMRG study. The parameters $t_{\parallel}$ and $J_{\parallel}$ take the same values as the ones in the SBMF study while $\tilde{J}_{\perp}=0.8J_{\parallel}$ is adopted in the DMRG study. To characterize the pairing symmetry and strength, we analyze the interlayer pairing correlation function $\Phi^{\perp}(r)$ and the intra-bottom-layer pairing correlation function $\Phi^{\parallel}_{b}(r)$. More details are provided in METHODS.
Fig. 4 (a) shows the pairing phase diagram with respect to $\delta$ ($=0,1/16,1/8$) and $\varepsilon$ ($\in[0,1.6t_{\parallel}]$). Fig. 4 (b) shows the absolute value of the intra-bottom-layer pairing correlation functions $|\Phi^{\parallel}_{b}(r)|$ under different electric fields $\varepsilon=0,0.4t_{\parallel},0.8t_{\parallel},1.2t_{\parallel},1.6t_{\parallel}$ for $\delta=0$, and the results for $\delta=1/16$ are presented in Fig. 4 (c). It turns out that $|\Phi^{\parallel}_{b}(r)|$ exhibits algebraic decay under an non-zero external electric field with the decaying power exponent to be $K_{\text{SC}}$, i.e. $|\Phi^{\parallel}_{b}(r)|\propto r^{-K_{\text{SC}}}$ for large enough $r$, implying the presence of pairing within the bottom layers. Fig. 4 (d) and (e) depict $|\Phi^{\parallel}_{b}(r)|$ for different $d_{x^{2}-y^{2}}$-electron-doping levels $\delta=0,1/16,1/8$ under $\varepsilon=0.4t_{\parallel}$ in (d) and $\varepsilon=0.8t_{\parallel}$ in (e). All the algebraic decay exponents $K_{\text{SC}}$ are provided accordingly.
The results indicate that (i) With the enhancement of the perpendicular electric field $\varepsilon$, and hence the $d_{x^{2}-y^{2}}$-electron-doping level $\delta$, the pairing symmetry changes from interlayer $s$-wave to intra-bottom-layer $d$-wave (The criterion of the pairing symmetry is provided in METHODS); (ii) For all the $d_{x^{2}-y^{2}}$-electron-doping levels $\delta$ tested, the enhancement of the perpendicular electric field $\varepsilon$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing; (iii) Under all the perpendicular electric field strengths $\varepsilon$ tested, the enhancement of the $d_{x^{2}-y^{2}}$-electron-doping level $\delta$ leads to a reduction of $K_{\text{SC}}$, suggesting the enhancement of the intra-bottom-layer pairing. From (ii) and (iii), it is clear that the enhancement of $\varepsilon$ and hence $\delta$ will significantly enhance the intra-bottom-layer pairing. These results are qualitatively consistent with those of the SBMF study. More results are given in the SI [146].
Figure 4: The DMRG results. (a) The $\delta-\varepsilon$ phase diagram of the ground state. The red region corresponds to the $s$-wave pairing and the blue region to the $d$-wave pairing. (b)-(c) The absolute value of the intra-bottom-layer pairing correlation functions $|\Phi^{\parallel}_{b}(r)|$ under different electric fields $\varepsilon=0,0.4t_{\parallel},0.8t_{\parallel},1.2t_{\parallel},1.6t_{\parallel}$ for $\delta=0$ in (b) and $\delta=1/16$ in (c). (d)-(e) $|\Phi^{\parallel}_{b}(r)|$ for different $d_{x^{2}-y^{2}}$-electron-doping levels $\delta=0,1/16,1/8$ under $\varepsilon=0.4t_{\parallel}$ in (d) and $\varepsilon=0.8t_{\parallel}$ in (e). The algebraic decay exponents $K_{\text{SC}}$ are written in the four figures as well. In (a-e), $\delta$ and $\varepsilon$ are set as independent variables, since their exact relationship is unclear.
I.2 The comprehensive two-orbital study
The above simplified single-orbital study has drawbacks: We cannot determine the relationship between the electron-doping of the $d_{x^{2}-y^{2}}$ orbitals and the electric field. In the SBMF study, we have to assume the ratio between the changes of the filling fractions of the two top-layer $E_{g}$ orbitals. In addition, we do not know how the neglected $3d_{z^{2}}$ orbital degree of freedom affects the pairing nature. To settle these puzzles, we conduct a comprehensive two-orbital model [107] to study with,
$H=$displaystyle-- $t_{\parallel}\sum_{\langle i,j\rangle,\mu}\left(c^{\dagger}_{i\mu x\sigma}c_{j\mu x\sigma}+\mathrm{h.c.}\right)-t_{\perp}\sum_{i}\left(c^{\dagger}_{itz\sigma}c_{ibz\sigma}+\mathrm{h.c.}\right)-t_{xz}\sum_{\langle i,j\rangle\mu}\left(c^{\dagger}_{i\mu x\sigma}c_{j\mu z\sigma}+(z\leftrightarrow x)+\mathrm{h.c.}\right)$ (2)
$+$ $J_{\parallel}\sum_{\langle i,j\rangle\mu}\bm{S}_{i\mu x}\cdot\bm{S}_{j\mu x}+J_{\perp}\sum_{i}\bm{S}_{itz}\cdot\bm{S}_{ibz}+\tilde{J}_{\perp}\sum_{i}\bm{S}_{itx}\cdot\bm{S}_{ibx}+\epsilon_{z}\sum_{i\mu\sigma}n_{i\mu z\sigma}+\epsilon_{x}\sum_{i\mu\sigma}n_{i\mu x\sigma}$
$+$ $\frac{\varepsilon}{2}\sum_{i\alpha\sigma}n_{it\alpha\sigma}-\frac{\varepsilon}{2}\sum_{i\alpha\sigma}n_{ib\alpha\sigma}.$
The operators $c_{i\mu\alpha\sigma}$, $n_{i\mu\alpha}$, $\bm{S}_{i\mu\alpha}$ take the same meanings as those in model (1) except for an extra index $\alpha=x/z$ labeling the orbital. Note that $\bm{S}_{i\mu\alpha}$ for each orbital is spin-$\frac{1}{2}$ operator. $\epsilon_{\alpha}$ denotes the on-site energy of the orbital $\alpha$. We adpot the tight-binding (TB) parameters reported in Ref. [141], i.e. $t_{\parallel}=0.445\text{ eV}$, $t_{xz}=0.221\text{ eV}$, $t_{\perp}=0.503\text{ eV}$ and $\epsilon_{x}-\epsilon_{z}=0.367\text{ eV}$. The superexchange interactions are obtained through $J\approx 4t^{2}/U$, with $U=10t_{\parallel}$. Finally $\varepsilon$ denotes the voltage between the two layers. More details are provided in METHODS.
[…]
The result shown in Fig. 5(b) for the comprehensive two-orbital study and that shown in Fig. 3(b) for the simplified one-orbital study look similar, except that in Fig. 5(b) the result is expressed as function of the directly controllable quantity $\varepsilon$. Actually, if we replace the $x$-axis of Fig. 5(b) by the calculated $n_{bx}=1-\delta_{bx}$, the resulting curve nearly coincides with Fig. 3(b), particularly in the large-$n_{bx}$ regime, see the SI [146]. The main reason for such similarity lies in that under strong $\bm{\varepsilon}$, the dominant superconducting pairing is the intra-bottom-layer $3d_{x^{2}-y^{2}}$-orbital pairing, which is insensitive to the $3d_{z^{2}}$ orbital. The main new information obtained in the two-orbital study lies in that the $3d_{z^{2}}$ orbital form interlayer $s$-wave pseudo-gap which is mixed with the intra-bottom-layer $d$-wave HTSC of the $3d_{x^{2}-y^{2}}$ orbital in the ratio of $1:\mathrm{i}$, as shown in Fig. 5(d). This state breaks time-reversal symmetry, although the experimentally detected superconducting gap is the standard $d$-wave gap of the $3d_{x^{2}-y^{2}}$ orbital. This intriguing result is left for experimental verification.
[…]
II Discussion
In conclusion, we propose that an imposed strong perpendicular electric field can drive HTSC with $T_{c}$ above the boiling point of liquid nitrogen in the single-bilayer film of La3Ni2O7 at AP. The reason lies in that under the strong electric field, the electrons in the layer with higher potential energy will flow to the layer with lower potential energy, to fill the $3d_{x^{2}-y^{2}}$ orbitals in the latter layer. With considerably enhanced filling fraction, the $3d_{x^{2}-y^{2}}$ electrons in that layer just mimic the cuprates, which form intralayer $d$-wave HTSC with strongly enhanced $T_{c}$. Our combined one-orbital and two-orbital studies consistently verify this idea.
Presently, while different groups have provided slightly different TB parameters for the La3Ni2O7 ultrathin film grown on SLAO substrate at AP, we have just adopted one set of these TB parameters to perform our calculations. However, the strong-coupling calculations performed here do not seriously rely on the accurate values of these parameters, because the main physics here is clear and simple. Actually, the well consistency between the result of the comprehensive two-orbital study and those of the simplified one-orbital studies with assuming different input conditions just verifies the robustness of our conclusion.
Moreover, we want to emphasize that the essence of introducing the perpendicular electric field is breaking the symmetry of the two layers by making their filling fractions different to each other. Actually, the filling fractions of different NiO layers in the La3Ni2O7 ultrathin film grown on the SLAO substrate may different from each other because of the existence of the substrate on one side of the film. This can be considered as effective electric field. Thus, our work provides a possible way to understand the high $T_{c}$ of the La3Ni2O7 ultrathin film grown on the SLAO substrate.
III Methods
III.1 The one-orbital model
The SBMF theory is used to solve the one-orbital model (1). In the SBMF approach, the superexchange terms are decomposed in $\chi-\Delta$ channel, e.g. $\bm{S}_{it}\cdot\bm{S}_{ib}=-\frac{3}{8}\left(\left\langle\chi^{\perp{\dagger}}\right\rangle\chi^{\perp}_{i}+\mathrm{h.c.}+\left\langle\Delta^{\perp{\dagger}}\right\rangle\Delta^{\perp}_{i}+\mathrm{h.c.}\right)$, $\chi$ and $\Delta$ represents hopping and pairing operators respectively. These MF parameters are further solved in a self-consistent manner. The specific steps can be referenced from prior work [79, 145, 107] and SI [146].
We also employ the DMRG method [147, 148] to solve the ground state of the Hamiltonian (1) as a comparison for the SBMF approach. The tensor libraries TensorKit [149] and FiniteMPS [150] provide an implementation of the required symmetry [151, 152]. We study the model on a $2\times 2\times L_{x}$ lattice with the open boundary conditions in the $x$ direction and choose $L_{x}=64$ for calculations. The matrix product state is constructed as shown in Fig. 6. We keep up to $D=12000$ $\mathrm{U}(1)_{\text{charge}}\times\mathrm{SU}(2)_{\text{spin}}$ multiplets in DMRG simulations and ensure the convergence accuracy of $10^{-6}$.
Figure 6: Illustration of the zigzag path in DMRG calculation.
The interlayer and intralayer singlet pairing operators take the form of
$\Delta^{\perp{\dagger}}_{i}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{it\uparrow}c^{{\dagger}}_{ib\downarrow}-c^{{\dagger}}_{it\downarrow}c^{{\dagger}}_{ib\uparrow}\right),$ (3)
$\Delta^{\parallel{\dagger}}_{i\mu}\equiv\Delta^{\bm{x}{\dagger}}_{i\mu}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{i\mu\uparrow}c^{{\dagger}}_{i+\bm{x},\mu\downarrow}-c^{{\dagger}}_{i\mu\downarrow}c^{{\dagger}}_{i+\bm{x},\mu\uparrow}\right),$
$\Delta^{\bm{y}{\dagger}}_{i\mu}=\frac{1}{\sqrt{2}}\left(c^{{\dagger}}_{i\mu\uparrow}c^{{\dagger}}_{i+\bm{y},\mu\downarrow}-c^{{\dagger}}_{i\mu\downarrow}c^{{\dagger}}_{i+\bm{y},\mu\uparrow}\right).$
Here, the subscripts $i+\bm{x}$($i+\bm{y}$) represent the NN site of the site $\bm{i}$ in the $x$($y$) direction. Fig. 7 shows how the singlet pairing operators are defined.
Figure 7: Illustration of the singlet pairing operators $\Delta^{\perp\dagger}_{i}$, $\Delta^{\parallel\dagger}_{i\mu}$ and $\Delta^{\bm{y}\dagger}_{i\mu}$.
The considered correlation functions are defined as follow
$\Phi^{\perp}(r)=\left\langle\Delta^{\perp{\dagger}}_{i}\Delta^{\perp}_{j}\right\rangle,$ (4)
$\Phi^{\parallel}_{\mu}(r)\equiv\Phi^{\bm{xx}}_{\mu}(r)=\left\langle\Delta_{i\mu}^{\parallel{\dagger}}\Delta^{\parallel}_{j\mu}\right\rangle,$
$\Phi^{\bm{xy}}_{\mu}(r)=\left\langle\Delta_{i\mu}^{\parallel{\dagger}}\Delta^{\bm{y}}_{j\mu}\right\rangle,$
where $r=|\bm{i}-\bm{j}|$ is the distance between the sites $i$ and $j$.
For a pairing channel whose absolute value of correlation function decays algebraically with distance, following the form $r^{-K_{\text{SC}}}$, the decay exponent $K_{\text{SC}}$ is associated with the Luttinger parameter specific to the channel. $K_{SC}<2$ signals a divergent superconducting susceptibility in that channel. The channel with the lowest $K_{\text{SC}}$ value is considered to dominate the pairing behavior.
The dominant pairing channel is related to pairing symmetry. For the case where interlayer pairing dominates, the pairing symmetry is restricted to $s$-wave pairing; while for the case where intralayer pairing in the bottom-layer dominates, we determine the pairing symmetry by the sign function $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\bm{xy}}_{b}(r)\right]$. If $\mathrm{sgn}\left[\Phi^{\parallel}_{b}(r)\Phi^{\bm{xy}}_{b}(r)\right]=-1$ holds for all $r$, the ground state can be identified as the $d$-wave pairing SC state. See SI [146] for more details on DMRG.
III.2 The two-orbital model
Here we provide more technique details for the SBMF study on the two-orbital model (2). The electron operator is decomposed as $c^{{\dagger}}_{i\mu\alpha\sigma}=f^{{\dagger}}_{i\mu\alpha\sigma}b_{i\mu\alpha}$, where $f$ is spinon operator and $b$ is holon operator. Since we have found that $T_{\mathrm{BEC}}\gg T_{\mathrm{pair}}$ in the considered $n_{bx}$ regime and $T_{\mathrm{pair}}$ is proportional to the zero-temperature spinon pairing gap, we can get the critical temperature of superconductivity only by calculating the ground-state spinon pairing gap. Thus we only consider the spinon Hamiltonian at zero temperature. The superexchange term is also decomposed in $\chi-\Delta$ channel. The spinon Hamiltonian is described as
$H_{\mathrm{spinon}}=$ $-t_{\parallel}\sum_{\langle i,j\rangle,\mu}\delta_{\mu x}\left(f^{\dagger}_{i\mu x\sigma}f_{j\mu x\sigma}+\mathrm{h.c.}\right)$ (5)
$-t_{xz}\sqrt{\delta_{tx}\delta_{tz}}\sum_{\langle i,j\rangle}\left(f^{\dagger}_{itx\sigma}f_{jtz\sigma}+f^{\dagger}_{itz\sigma}f_{jtx\sigma}+\mathrm{h.c.}\right)$
$-\frac{3}{8}J_{\parallel}\sum_{\langle i,j\rangle\mu}\left(\chi_{ij,\mu x}^{\dagger}\left\langle\chi_{\mu x}\right\rangle+\mathrm{h.c.}-\left\langle\chi_{\mu x}^{\dagger}\right\rangle\left\langle\chi_{\mu x}\right\rangle\right)$
$-\frac{3}{8}J_{\parallel}\sum_{\langle i,j\rangle\mu}\left(\Delta_{ij,\mu x}^{\dagger}\left\langle\Delta_{\mu x}\right\rangle+\mathrm{h.c.}-\left\langle\Delta_{\mu x}^{\dagger}\right\rangle\left\langle\Delta_{\mu x}\right\rangle\right)$
$-\frac{3}{8}J_{\perp}\sum_{i}\left(\chi^{\perp{\dagger}}_{iz}\left\langle\chi^{\perp}_{z}\right\rangle+\mathrm{h.c.}-\left\langle\chi^{\perp{\dagger}}_{z}\right\rangle\left\langle\chi^{\perp}_{z}\right\rangle\right)$
$-\frac{3}{8}J_{\perp}\sum_{i}\left(\Delta^{\perp{\dagger}}_{iz}\left\langle\Delta^{\perp}_{z}\right\rangle+\mathrm{h.c.}-\left\langle\Delta^{\perp{\dagger}}_{z}\right\rangle\left\langle\Delta^{\perp}_{z}\right\rangle\right)$
$-\frac{3}{8}\tilde{J}_{\perp}\sum_{i}\left(\Delta^{\perp{\dagger}}_{ix}\langle\Delta^{\perp}_{x}\rangle+\mathrm{h.c.}-\langle\Delta^{\perp{\dagger}}_{x}\rangle\langle\Delta^{\perp}_{x}\rangle\right)$
$+\sum_{i\mu\alpha\sigma}\epsilon_{\alpha}n_{i\mu\alpha\sigma}+\frac{\varepsilon}{2}\sum_{i\alpha\sigma}n_{it\alpha\sigma}-\frac{\varepsilon}{2}\sum_{i\alpha\sigma}n_{ib\alpha\sigma}.$
where $\delta_{\mu\alpha}=\left\langle{}b_{i\mu\alpha}b^{{\dagger}}_{j\mu\alpha}\right\rangle$ since holon condense at zero temperature. Under the electric field, we have $\delta_{bz}=0$ and $\delta_{tz}$, $\delta_{tx}$ and $\delta_{bx}$ are solved in a self-consistent manner by adjustment to onsite energies $\epsilon_{\alpha}$ (See SI [146] for more details). The mean-field order parameters are represented by
$\chi_{ij,\mu x}=\sum_{\sigma}f_{i\mu x\sigma}^{\dagger}f_{j\mu x\sigma},$ (6)
$\chi^{\perp{\dagger}}_{iz}=\sum_{\sigma}f^{\dagger}_{izt\sigma}f_{izb\sigma},$
$\Delta^{\dagger}_{ij,\mu\alpha}=f^{\dagger}_{i\mu\alpha\uparrow}f^{\dagger}_{j\mu\alpha\downarrow}-f^{\dagger}_{i\mu\alpha\downarrow}f^{\dagger}_{j\mu\alpha\uparrow},$
$\Delta^{\perp{\dagger}}_{i\alpha}=f^{\dagger}_{it\alpha\uparrow}f^{\dagger}_{ib\alpha\downarrow}-f^{\dagger}_{ib\alpha\downarrow}f^{\dagger}_{it\alpha\uparrow},$
and
$\chi_{\mu x}=\frac{1}{2N}\sum_{\langle i,j\rangle}\chi_{ij,\mu x},\ \chi^{\perp}_{z}=\frac{1}{N}\sum_{i}\chi^{\perp}_{iz},$ (7)
$\Delta^{\bm{x}}_{\mu x}=\frac{1}{2N}\sum_{\langle i,j\rangle}\Delta_{ij,\mu x},\ \Delta^{\perp}_{\alpha}=\frac{1}{N}\sum_{i}\Delta^{\perp}_{i\alpha}.$
Notably, the spin-exchange $\tilde{J}_{\perp}$ of the Hamiltonian in Eq. (2) doesn’t produce a hopping term $\chi_{x}^{\perp}$ in Eq. (5), which is the feature of such a bilayer system. Without interlayer hopping, a small interlayer spin-exchange $J_{\perp}$ leads to $\langle\chi^{\perp}\rangle\approx 0$.
Consequently, the $3d_{z^{2}}$ orbital only participate in the interlayer pairing. However, this pairing is not SC as the corresponding SC order parameter goes to zero in the SBMF theory due to $\delta_{bz}=0$. The SC is carried by the $3d_{x^{2}-y^{2}}$ orbitals, which can form both intralayer and interlayer pairing. The superconducting $T_{c}$ scales with the ground state gap amplitude of the $3d_{x^{2}-y^{2}}$ orbitals via the BCS relation exhibited in Fig. 3(b).
The expectation value of the mean-field order parameters are obtained by numerically solving the following self-consistent equations
$\delta_{\mu\alpha}=1-\frac{1}{N}\sum_{k}\left(\langle f^{\dagger}_{k\mu\alpha\uparrow}f_{k\mu\alpha\uparrow}\rangle+\langle f^{\dagger}_{-k\mu\alpha\downarrow}f_{-k\mu\alpha\downarrow}\rangle\right),$ (8)
$\delta_{tz}=0,\ \sum_{\mu\alpha}\delta_{\mu\alpha}=1,$
$\langle\chi_{\mu x}\rangle=\frac{1}{N}\sum_{k}\epsilon(\bm{k})\left(\langle f^{\dagger}_{k\mu x\uparrow}f_{k\mu x\uparrow}\rangle+\langle f^{\dagger}_{-k\mu x\downarrow}f_{-k\mu x\downarrow}\rangle\right),$
$\langle\chi^{\perp}_{z}\rangle=\frac{1}{N}\sum_{k}\left(\langle f^{\dagger}_{ktz\uparrow}f_{kbz\uparrow}\rangle+\langle f^{\dagger}_{-ktz\downarrow}f_{-kbz\downarrow}\rangle\right),$
$\langle\Delta^{\bm{x}}_{\mu x}\rangle^{*}=\frac{1}{N}\sum_{k}2\cos{(k_{x})}\langle f^{\dagger}_{k\mu x\uparrow}f^{\dagger}_{-k\mu x\downarrow}\rangle,$
$\langle\Delta^{\perp}_{\alpha}\rangle^{*}=\frac{2}{N}\sum_{k}\langle f^{\dagger}_{kt\alpha\uparrow}f^{\dagger}_{-kb\alpha\downarrow}\rangle,$
where $\epsilon(\bm{k})=\frac{\cos{(k_{x})}+\cos{(k_{y})}}{2}$.
IV Data availability
All data are displayed in the main text and the supplementary information.
V Code availability
The code that supports the plots within this paper is available from the corresponding author upon request.
VI Acknowledgement
We are grateful to the stimulating discussions with Chen Lu. F. Y. and C. W are supported by the National Natural Science Foundation of China (NSFC) under the Grant No. 12234016, and also supported by the NSFC under the Grant Nos. 12074031 and 12174317, respectively. D. X. Y. is supported by NKRDPC-2022YFA1402802, NSFC-92165204, and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).
VII Author contributions
F. Yang proposed the main idea and supervised the study. Z.-Y. Shao performed the SBMF study. J.-H. Ji performed the DMRG study. F. Yang, Z.-Y. Shao and J.-H. Ji wrote the paper. All authors significantly contributed to the data analysis and discussion. .